# Entropy Rectifying Guidance for Diffusion and Flow Models

### v1

@@ 3 Guidance via Entropy Rectification
Similarly to previous approaches (Ho and Salimans, 2021; Hong, 2024; Karras et al., 2024), our approach is based on contrasting the denoising estimate with a less powerful one. In particular, using the Hopfield energy interpretation of attention in Equation 4, we manipulate the energy landscape of the attention layer, by rectifying the entropy of the associations in the attention operation. The modified attention layers lead to lower quality predictions, which are used as the contrasting term for guidance. This approach does not require a second model and can be used for both conditional and unconditional sampling. We refer to our approach as Entropy Rectifying Guidance (ERG).
@@ 3.1 Manipulating the energy landscape
The temperature rescaling parameter $\tau$ controls the sharpness of the softmax attention, and $\alpha$ the relative importance of the similarity between the state matching term compared to the norm of the state patterns. Temperature rescaling with $\tau$ is similar to the Gaussian blurring of the attention maps introduced in SEG  (Hong, 2024), but allows for non-local smoothing of the attention layers. Additionally, the view of the attention layer as a CCCP update of the energy function, allows for consideration of different methods to minimize the energy landscape. For instance, taking $K$ gradient descent steps with step size $\gamma$, as illustrated in Algorithm 1. Using different settings of the hyperparameters $\alpha,\gamma,$ and $\tau$ allows us to manipulate the energy landscape modeling the attention operation, and obtain noise estimates that deviate from the trained model. We expect these to be weaker estimates compared to those provided by the model, as it was trained with standard attention layers, i.e. with $\alpha=\gamma=\tau=1$. When applying this rectification mechanism to the denoiser model, we refer to the method as image-ERG, or I-ERG for short. In particular, we apply it to the negative/unconditional prediction part of the classifier-free guidance, only on certain layers of the network that will be defined subsequently. Additionally, we impose a kickoff threshold $\kappa$ on the time steps in which guidance is applied.
@@ 3.2 Manipulating the energy of the text encoder
Besides the image denoising model, we can also manipulate the energy landscape of attention-based text encoders to obtain a weak version of the conditional embeddings. Let $c$ be the text prompt that is used as conditioning. The text tokens are obtained by feeding the prompt to the text encoder: $\textrm{Enc}(c)\in\mathbb{R}^{d_{t}}$, such as Llama (Grattafiori et al., 2024) or T5 (Chung et al., 2022). Text tokens are then fed to the denoiser model through cross-attention layers.
@@ 3.3 Guidance update
where $D$ is the learned denoiser model with parameters $\Theta$ that are omitted for simplicity, $D^{\boldsymbol{\xi}}$ is the denoiser model where the attention layers have been replaced by the modified version presented in Equation 9, and $\Theta_{\xi}$ the set of hyperparameters introduced by our method. $\boldsymbol{\phi}_{c}$ are hidden states of the text encoder while $\boldsymbol{\phi}_{c}^{\tau}$ is the hidden states of the text encoder obtained with a modified energy landscape in the attention layers. Compared to the standard CFG, the main differences are that we replace the unconditional text embeddings with conditional embeddings obtained with the entropy-rectified attention mechanism (C-ERG). Similarly, the denoiser model used for the negative/unconditional predictions is also obtained with a rectified attention mechanism (I-ERG). Finally, the denoiser level updates are only applied after a certain point during sampling in order to not overly penalize the negative components of the ERG update at the start of sampling.
@@ 4.1 Experimental setup
Datasets and architectures.  We experiment with class-conditional and text-to-image models. Both models are trained at $512\times 512$ image resolution using the rectified flow-matching paradigm (Lipman et al., 2023; Esser et al., 2024). We train class-conditional models on a faceblurred version of ImageNet (Deng et al., 2009). For the class-conditional setting, we experiment with the XL/2 variant of the DiT architecture (Peebles and Xie, 2023) which is composed of 28 attention blocks with hidden dimension of $1152$, resulting in $790$M parameters; in this setting we evaluate models at both $256$ and $512$ resolutions. For the text-to-image model, we use an architecture similar to MMDiT (Esser et al., 2024), and train on a mix of a proprietary dataset of 320M text-image pairs and YFCC100M, where all faces in YFCC100M have been blurred. Similarly to MMDiT (Esser et al., 2024), the model uses a mix of different text encoders: Llama3-8b (Grattafiori et al., 2024) and Flan-T5-XL (Chung et al., 2022). During training, each of the text encoders is disabled with a probability of $\sqrt{0.1}$, so that the probability of both encoders being disabled is around $10\%$. We enable both text encoders during inference time. The architecture of the model is made of $38$ blocks with a hidden dimension of $1536$, resulting in approx $1.9$B parameters. For both datasets, we recaption the images using both Florence-2 Large (Xiao et al., 2024) to obtain medium-length captions and PaliGemma-3b (Beyer et al., 2024) for shorter COCO-style captions. Additional techniques to improve training efficiency of this model, such as conditioning mechanisms and pre-training strategies, were adopted from Berrada et al. (2024). For the class-conditional model we use the asymmetric autoencoder of Zhu et al. (2023), while for the larger text-to-image model we use the SD3 autoencoder which has latent representation with 16 channels.
Metrics.  We consider metrics for quality, diversity, and consistency (Astolfi et al., 2024). We measure sample quality with FID (Heusel et al., 2017), precision (Kynkäänniemi et al., 2019) and density (Naeem et al., 2020); sample diversity is measured with recall (Kynkäänniemi et al., 2019), coverage (Naeem et al., 2020) and FID; and input consistency is measured with CLIPScore (Hessel et al., 2021) and VQAScore (Lin et al., 2024).
@@
For evaluation of text-to-image and unconditional generation, we use the 40k COCO’14 validation image-caption pairs. For the class-conditional models, we sample 50 images for each of the 1,000 ImageNet classes and use the ImageNet validation set as a reference. All evaluated models are sampled using the Euler method with $50$ sampling steps. We use the EvalGIM (Hall et al., 2024) library for all evaluations.
@@
Table 1: Comparison of our Entropy Rectifying Guidance (ERG) with other guidance approaches. Conducted for text-to-image generation using the COCO’14 validation data. Scores obtained by the hyper-parameter setting achieving the optimal mean rank across the pool of combinations. We compare ERG to other state-of-the-art guidance approaches and mark the best result in each column in bold in the top part of the table. In the bottom part of the table we evaluate combinations of ERG with APG and CADS, and bold results when they surpass the results reported in the upper part of the table.
@@
Guidance Metric FID ()($\downarrow$)( ↓ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ ) Density ()($\uparrow$)( ↑ ) Coverage ()($\uparrow$)( ↑ ) CLIPScore ()($\uparrow$)( ↑ ) VQAScore ()($\uparrow$)( ↑ ) NFE ()($\downarrow$)( ↓ )
CFG (Ho and Salimans, 2021) $12.81$ $65.71$ $43.95$ $98.24$ $71.12$ $26.45$ $70.15$ $2$
APG (Sadat et al., 2025) 11.8811.88{11.88}11.88 $66.34$ $45.15$ $104.07$ $73.06$ $26.54$ $72.47$ $2$
CADS (Sadat et al., 2024) $11.93$ $66.42$ $45.75$ $101.01$ $72.99$ $26.76$ $73.36$ $2$
PAG* (Ahn et al., 2024) $12.75$ $66.42$ $43.94$ $107.21$ $72.20$ $26.80$ $73.32$ $2$
SAG* Hong et al. (2023) ${\bf 11.68}$ $64.97$ ${\bf 47.98}$ $103.58$ $72.74$ $26.81$ $72.16$ $2$
SEG* (Hong, 2024) $16.87$ $61.22$ $36.88$ $87.77$ $61.91$ ${\bf 26.86}$ $73.59$ $3$
AutoGuidance (Karras et al., 2024) $16.62$ $61.48$ $34.60$ $87.02$ $62.75$ $26.59$ $73.53$ $2$
ERG (ours) $13.62$ $\bf 70.92$ $41.43$ $\bf 120.25$ $\bf 73.21$ ${\bf 26.86}$ ${\bf 73.96}$ $2$
ERG (ours) + APG ${\bf 11.37}$ $69.50$ ${\bf 50.25}$ $115.08$ ${\bf 80.50}$ $26.74$ $73.55$ $2$
ERG (ours) + CADS $12.87$ ${\bf 72.72}$ $38.89$ ${\bf 128.54}$ $76.23$ $26.75$ $73.45$ $2$
@@
Baselines.  In addition to the standard classifier-free guidance, we compare our method with Condition-Annealed Diffusion Sampler (CADS) (Sadat et al., 2024), Adaptive Projected Guidance (APG) (Sadat et al., 2025), Smooth Energy Guidance (SEG) (Hong, 2024) and Auto-Guidance (Karras et al., 2024). For APG, we follow the recommendations from the paper and set $\gamma_{\text{APG}}=-0.5$, $\eta_{\text{APG}}=0.0$, $r_{\text{APG}}=5.0$. For CADS, we perform a grid search over $\tau_{1}^{\text{CADS}}\in[0.6,0.8],\tau_{2}^{\text{CADS}}\in[0.8,1.0],s^{\text{CADS}}\in[0.25,1.0],\psi^{\text{CADS}}=1.0$. To ensure a fair comparison, we select the best performing guidance strength for baselines as well as our method using the rank-scoring algorithm defined in the appendix, see Section A.1 for details. Note that this is different from reporting the optimal score achieved for each metric, as such a report might not correspond to any particular run because of inherent trade-offs between the different facets of the generations.
Since SAG, PAG and SEG were developed specifically for the UNet architecture, we adapt these method for diffusion transformers (Peebles and Xie, 2023) by applying the method in the attention layers of the middle blocks of the transformer; we refer to these methods with an asterisk superscript *. Additional details on the choice of layers are provided in Section B.4. For AutoGuidance, we follow the recommendations from the paper and use an earlier checkpoint of the same model, at approximately $1/16$-th of the training, as the weaker model.
@@ 4.2 Main experimental results
Figure 2: Pareto fronts on consistency-diversity-quality for text-to-image generation. Metrics computed on COCO using ERG + APG, points with a $\times$ sign correspond to APG. We sweep over different guidance scales (each marked with a different color), and hyper-parameters $\alpha,\gamma,\tau$ for ERG. Dashed lines trace the Pareto fronts for each plot. We measure consistency with VQAScore, quality with density and diversity with coverage.
Text-to-image generation.  In Table 1, we compare our Entropy Rectifying Guidance with different guidance mechanisms from the literature on the text-to-image generation task. Our proposed ERG approach demonstrates competitive performance across various metrics for text-to-image generation on the COCO’14 validation set, outperforming baselines such as CFG, SAG, PAG and SEG* in most metrics. Specifically, ERG considerably boosts image quality as reported in quality metrics (e.g., +22 for density and +5 for precision when compared to CFG). Additionally, ERG combined with APG achieves the overall best FID, recall and coverage, and second-best overall precision and density. Notably, our approach requires only two function evaluations (NFE) per inference step, which is comparable to other methods except SEG which requires three. These results suggest that our ERG approach is successfully able to boost the generation quality of the model in all three facets of generation (quality, diversity, and consistency).
Following Astolfi et al. (2024), we plot Pareto fronts for quality (measured by density), diversity (Coverage), and consistency (VQAScore) metrics in Figure 2 when using ERG +APG. We find all the points belonging to the Pareto fronts correspond to ERG +APG with APG. ERG +APG is able to improve all three facets of the generations w.r.t. APG, and provides significant boosts for quality and consistency. This can also be observed in the qualitative examples in Figure 1, where images sampled using our method show better visual quality than the baseline with classifier-free guidance.
@@
Figure 3: Unconditional generation results. Compared to not using guidance, our ERG generates more realistic and detailed images and more coherent structure. Images obtained from T2I model at $512$ with empty prompt inputs.
Table 2: Unconditional generation. Comparing no guidance, SEG, and I-ERG sampling from the text-to-image model.
FID Precision Recall Density Coverage
No guidance $101.50$ $13.50$ $38.22$ $8.99$ $3.63$
SAG (Hong et al., 2023) $39.25$ $43.45$ $55.00$ $46.22$ $30.70$
PAG* (Ahn et al., 2024) $41.50$ $42.91$ ${\bf 62.86}$ $45.65$ $30.51$
SEG* (Hong, 2024) $37.75$ $48.99$ $50.10$ $55.56$ $34.79$
AutoGuidance (Karras et al., 2024) $39.50$ $49.97$ $42.09$ $48.26$ $34.71$
ERG (ours) ${\bf 36.25}$ ${\bf 68.87}$ $57.15$ ${\bf 55.84}$ ${\bf 51.59}$
Figure 4: Temperature rescaling in the conditioning (C-ERG) for text-to-image generation. We vary the attention temperatures $\tau_{c}$ on the text encoder. Each curve represents one setting of the guidance strength $w$. The left-most point on each curve represents the result for standard CFG with the at weight.
Table 3: Class-conditional generation. Comparison of ERG with other guidance methods for models trained for 256 and 512 resolution generation.
Res. FID Precision Recall Density Coverage
CFG (Ho and Salimans, 2021) $256$ ${\bf 3.67}$ $76.63$ $55.70$ $127.03$ $85.81$
PAG* (Ahn et al., 2024) $5.31$ $72.38$ $52.55$ $111.94$ $81.04$
SAG (Hong et al., 2023) $3.78$ $75.56$ $\bf 58.68$ $131.89$ $86.35$
SEG* (Hong, 2024) $6.15$ $77.97$ $48.10$ $132.22$ $84.51$
I-ERG (ours) ${\bf 3.67}$ ${\bf 80.51}$ $54.03$ ${\bf 141.96}$ ${\bf 86.72}$
CFG (Ho and Salimans, 2021) $512$ $5.65$ $81.39$ $49.06$ $146.97$ ${\bf 86.70}$
PAG* (Ahn et al., 2024) $4.65$ $78.04$ $52.44$ $134.49$ $86.50$
SAG (Hong et al., 2023) $4.81$ $74.90$ $52.84$ $120.09$ $83.91$
SEG* (Hong, 2024) $6.59$ $82.88$ $33.25$ $160.11$ $81.85$
I-ERG (ours) ${\bf 4.56}$ ${\bf 84.39}$ ${\bf 52.92}$ ${\bf 163.63}$ $86.13$
Unconditional generation.  In the unconditional model sampling experiment, we compare our Entropy Rectifying Guidance with sampling without guidance and using applicable methods: AutoGuidance, SEG*, PAG* and SAG*. We provide quantitative evaluation results in Table 2. We find that ERG outperforms all other methods, except for Recall which is higher for the PAG* method (62.86 vs. 57.15 for ERG), with significant boosts in FID, precision, and coverage over other methods. The qualitative examples in Figure 3 clearly show artifacts in terms of structural coherence in all the objects present in the generated images when not using guidance, which disappear when using ERG.
Class-conditional generation.  For class-conditional generation, we observe similar trends as the text-to-image setup in Table 3. Mainly, ERG boosts quality related metrics while maitaining similar performance in terms of diversity metrics (Recall and Coverage). At 256 resolution, we find large improvements in density, and smaller improvements in coverage and precision. FID remains similar to CFG but is better compared to all other methods tested. At $512$ resolution, ERG is best across all metrics, except for coverage where ERG is slightly behind CFG and SEG*.

### v2

@@ 3 Guidance via Entropy Rectification
Similarly to previous approaches (Ho and Salimans, 2021; Hong, 2024; Karras et al., 2024), our approach is based on contrasting the denoising estimate with a less powerful one. In particular, using the Hopfield energy interpretation of attention in Section˜2.3, we manipulate the energy landscape of the attention layer, by rectifying the entropy of the associations in the attention operation. The modified attention layers lead to lower quality predictions, which are used as the contrasting term for guidance. This approach does not require a second model and can be used for both conditional and unconditional sampling. We refer to our approach as Entropy Rectifying Guidance (ERG).
4 Guidance via entropy rectification
Similarly to previous approaches (Ho and Salimans, 2021; Hong, 2024; Karras et al., 2024), our approach is based on contrasting the (conditional) denoising estimate with a less powerful one. In particular, using the Hopfield energy interpretation of attention (Ramsauer et al., 2021), see Section˜2.3, we manipulate the energy landscape of the attention layer, by rectifying the entropy of the associations in the attention operation. The modified attention layers lead to lower quality predictions, which are used as the contrasting term for guidance. This approach does not require a second model, and can be used for both conditional and unconditional sampling. We refer to our approach as Entropy Rectifying Guidance (ERG).
4.1 Manipulating the energy landscape
$E(\boldsymbol{\xi};{\bf X})=\frac{1}{2}\boldsymbol{\xi}^{\top}\boldsymbol{\xi}-\alpha\cdot\text{LogSumExp}\left({\bf X}^{\top}\boldsymbol{\xi},\tau\cdot\beta\right),$ (1)
Algorithm 1 Multi-step Entropy Rectifying Guidance.
$K\in\mathbb{N}$ number of gradient update steps.
$\gamma>0$ step size.
$\alpha\in\mathbb{R}$ State pattern matching weight.
$\tau>0$ attention temperature.
${\bf K},{\bf V}\in\mathbb{R}^{d\times N}$ keys and values.
${\bf Q}\in\mathbb{R}^{d\times S}$ queries.
$k\leftarrow 0$
while $k<K$ do
  ${\bf Q}\leftarrow{\bf Q}-\gamma\left({\bf Q}-\alpha\cdot\text{softmax}\left(\tau\cdot\beta{\bf Q}{\bf K}^{\top}\right){\bf V}\right)$ 11footnotemark: 1
  $k\leftarrow k+1$
end while
where $\beta=\frac{1}{\sqrt{d}}$ is the default temperature of the attention attention update. The temperature rescaling parameter $\tau$ controls the sharpness of the softmax attention, and $\alpha$ the relative importance of the similarity between the state matching term compared to the norm of the state patterns. Temperature rescaling with $\tau$ is similar to the Gaussian blurring of the attention maps introduced in SEG (Hong, 2024), but allows for non-local smoothing of the attention maps. Additionally, the view of the attention layer as a CCCP update of the energy function, allows for consideration of different methods to minimize the energy landscape. For instance, taking $K$ gradient descent steps with step size $\gamma$, as illustrated in Algorithm˜2. Using different settings of the hyperparameters $\alpha,\gamma,$ $\tau$ and $K$ allows us to manipulate the attention operation, and obtain noise estimates that deviate from the trained model. We expect these to be weaker estimates compared to those provided by the model, as it was trained with standard attention layers, i.e. with $\alpha=\gamma=\tau=K=1$. When applying this rectification mechanism to the denoiser model, we refer to the method as image-ERG, or I-ERG for short. In particular, we apply it to the negative/unconditional prediction part of the classifier-free guidance, and only on certain layers of the network that we will identify in our experiments. Additionally, we impose a kickoff threshold $\kappa$ on the time steps in which guidance is applied.
4.2 Manipulating the energy of the text encoder
Besides the image denoising model, we can also manipulate the energy landscape of attention-based text encoders to obtain a weak version of the conditional embeddings. Let $c$ be the text prompt that is used as conditioning. The text tokens are obtained by feeding the prompt to the text encoder: $\textrm{Enc}(c)\in\mathbb{R}^{d_{t}}$, such as Llama (Grattafiori et al., 2024) or T5 (Chung et al., 2024a). Text tokens are then fed to the denoiser model through cross-attention layers.
To obtain a contrasting signal for guidance, we manipulate the energy landscape in the self-attention layers of the text encoder following Algorithm 2. More specifically, for every self-attention layer in the text encoder, we introduce a temperature hyperparameter in the softmax function. This enables us to change the strength by which keys and queries are being matched, resulting in a modified prompt embedding at the output. For the remainder of the manuscript, we refer to this method condition-ERG, or C-ERG for short. For simplicity reasons, for the text encoder, we only consider changing the temperature but not the step size $\gamma$, pattern matching weight $\alpha$, and number of update steps $K$.
4.3 Guidance update
where $D$ is the learned denoiser model with parameters $\Theta$ that are omitted for simplicity, $D^{\boldsymbol{\xi}}$ is the denoiser model where the attention layers have been replaced by the modified version presented in Equation˜3, and $\Theta_{\xi}=\{\alpha,\gamma,\tau,K\}$ the set of hyperparameters introduced by ERG. We use $\boldsymbol{\phi}_{c}$ to denote the prompt embedding produced by the text encoder, while $\boldsymbol{\phi}_{c}^{\tau}$ denotes the embedding obtained with the modified attention layers.
Compared to standard CFG, the main differences are that (i) we replace the unconditional text embeddings with conditional embeddings obtained with the entropy-rectified attention mechanism (C-ERG), and (ii) the denoiser model for the negative/unconditional predictions also uses modified attention layers (I-ERG). Finally, the changes to the image denoiser are only applied after a certain point during sampling in order to not overly penalize the negative components of the ERG update at the start of sampling. For the text-encoder we apply the temperature scaling throughout the denoising process, so that at any stage we obtain a noise prediction that can be contrasted with the vanilla denoising signal.
Note that our approach can be combined with other approaches, e.g., CADS (Sadat et al., 2024) and APG (Sadat et al., 2025). We will explore such combinations in our experiments.
@@ 4.4 Manipulating the energy landscape
The temperature rescaling parameter $\tau$ controls the sharpness of the softmax attention, and $\alpha$ the relative importance of the similarity between the state matching term compared to the norm of the state patterns. Temperature rescaling with $\tau$ is similar to the Gaussian blurring of the attention maps introduced in SEG (Hong, 2024), but allows for non-local smoothing of the attention layers. Additionally, the view of the attention layer as a CCCP update of the energy function, allows for consideration of different methods to minimize the energy landscape. For instance, taking $K$ gradient descent steps with step size $\gamma$, as illustrated in Algorithm˜2. Using different settings of the hyperparameters $\alpha,\gamma,$ and $\tau$ allows us to manipulate the energy landscape modeling the attention operation, and obtain noise estimates that deviate from the trained model. We expect these to be weaker estimates compared to those provided by the model, as it was trained with standard attention layers, i.e. with $\alpha=\gamma=\tau=1$. When applying this rectification mechanism to the denoiser model, we refer to the method as image-ERG, or I-ERG for short. In particular, we apply it to the negative/unconditional prediction part of the classifier-free guidance, only on certain layers of the network that will be defined subsequently. Additionally, we impose a kickoff threshold $\kappa$ on the time steps in which guidance is applied.
@@ 4.5 Manipulating the energy of the text encoder
Besides the image denoising model, we can also manipulate the energy landscape of attention-based text encoders to obtain a weak version of the conditional embeddings. Let $c$ be the text prompt that is used as conditioning. The text tokens are obtained by feeding the prompt to the text encoder: $\textrm{Enc}(c)\in\mathbb{R}^{d_{t}}$, such as Llama (Grattafiori et al., 2024) or T5 (Chung et al., 2024a). Text tokens are then fed to the denoiser model through cross-attention layers.
@@ 4.6 Guidance update
where $D$ is the learned denoiser model with parameters $\Theta$ that are omitted for simplicity, $D^{\boldsymbol{\xi}}$ is the denoiser model where the attention layers have been replaced by the modified version presented in Equation˜3, and $\Theta_{\xi}$ the set of hyperparameters introduced by our method. $\boldsymbol{\phi}_{c}$ are hidden states of the text encoder while $\boldsymbol{\phi}_{c}^{\tau}$ is the hidden states of the text encoder obtained with a modified energy landscape in the attention layers. Compared to the standard CFG, the main differences are that we replace the unconditional text embeddings with conditional embeddings obtained with the entropy-rectified attention mechanism (C-ERG). Similarly, the denoiser model used for the negative/unconditional predictions is also obtained with a rectified attention mechanism (I-ERG). Finally, the denoiser level updates are only applied after a certain point during sampling in order to not overly penalize the negative components of the ERG update at the start of sampling.
@@ 5.1 Experimental setup
Datasets and architectures.  We experiment with class-conditional and text-to-image models trained using rectified flow-matching (Lipman et al., 2023; Esser et al., 2024). We use a face-blurred version of ImageNet (Deng et al., 2009) to train class-conditional models at 256 and 512 resolution based on the XL/2 variant of the DiT architecture (Peebles and Xie, 2023), which is composed of 28 attention blocks with hidden dimension of $1152$, resulting in 790790M parameters. For the text-to-image model, we use an architecture similar to MMDiT (Esser et al., 2024), and train a 512 resolution model on a mix of a proprietary dataset of 320M text-image pairs and YFCC100M (Thomee et al., 2016), where all faces in YFCC100M have been blurred. Similar to MMDiT (Esser et al., 2024), the model uses a mix of different text encoders: Llama3-8B (Grattafiori et al., 2024) and Flan-T5-XL (Chung et al., 2024a). During training, each of the text encoders is disabled with a probability of $\sqrt{0.1}$, so that the probability of both encoders being disabled is around $10\%$. We enable both text encoders during inference time for text-to-image generation, and disable both text-encoders for unconditional image generation experiments. The architecture of the model is made of $38$ blocks with a hidden dimension of $1,536$, resulting in approx 1.91.9B parameters. For both datasets, we recaption the images using both Florence-2 Large (Xiao et al., 2024) to obtain medium-length captions and PaliGemma-3B (Beyer et al., 2024) for shorter COCO-style captions. Additional techniques to improve training efficiency of this model, such as conditioning mechanisms and pre-training strategies, were adopted from Berrada et al. (2024). For the class-conditional model we use the asymmetric autoencoder of Zhu et al. (2023), while for the larger text-to-image model we use the SD3 autoencoder (Esser et al., 2024).
@@
Metrics.  We consider metrics for quality, diversity, and consistency (Astolfi et al., 2024). We measure sample quality with FID (Heusel et al., 2017) and density (Naeem et al., 2020); sample diversity is measured with coverage (Naeem et al., 2020) and FID; and prompt consistency is measured with CLIPScore (Hessel et al., 2021) and VQAScore (Lin et al., 2024). For evaluation of text-to-image and unconditional generation, we use the 40k COCO’14 validation image-caption pairs. For the class-conditional models, we sample 50 images for each of the 1,000 ImageNet classes and use the ImageNet validation set as a reference. All evaluated models are sampled using the Euler method with $50$ sampling steps. We use the EvalGIM (Hall et al., 2024) library for all evaluations.
@@
Table 1: Comparison of ERG with other guidance approaches for text-to-image generation. We compare ERG to other state-of-the-art guidance approaches and mark the best result in each column in bold in the top part of the table. In the bottom part of the table we evaluate combinations of ERG with APG and CADS, and bold results when they surpass the results in the upper part of the table.
@@
Guidance Metric FID $(\downarrow)$ Density $(\uparrow)$ Coverage $(\uparrow)$ CLIPScore $(\uparrow)$ VQAScore $(\uparrow)$ NFE $(\downarrow)$
CFG (Ho and Salimans, 2021) $12.81$ $98.24$ $71.12$ $26.45$ $70.15$ 22
APG (Sadat et al., 2025) 11.88{11.88} $104.07$ $73.06$ $26.54$ $72.47$ 22
CADS (Sadat et al., 2024) $11.93$ $101.01$ $72.99$ $26.76$ $73.36$ 22
PAG* (Ahn et al., 2024) $12.75$ $107.21$ $72.20$ $26.80$ $73.32$ 22
SAG* Hong et al. (2023) $\bf 11.68$} $103.58$ $72.74$ $26.81$ $72.16$ 22
SEG* (Hong, 2024) $16.87$ $87.77$ $61.91$ $\bf 26.86$} $73.59$ 33
AutoGuidance (Karras et al., 2024) $16.62$ $87.02$ $62.75$ $26.59$ $73.53$ 22
ERG (ours) $13.62$ $\bf 120.25$ $\bf 73.21$ $\bf 26.86$} $\bf 73.96$} 22
ERG (ours) + APG $\bf 11.37$} $115.08$ $\bf 80.50$} $26.74$ $73.55$ 22
ERG (ours) + CADS $12.87$ $\bf 128.54$} $76.23$ $26.75$ $73.45$ 22
@@
Baselines.  In addition to the standard classifier-free guidance, we compare our method to several recent state-of-the-art guidance techniques: Condition-Annealed Diffusion Sampler (CADS) (Sadat et al., 2024), Adaptive Projected Guidance (APG) (Sadat et al., 2025), Smooth Energy Guidance (SEG) (Hong, 2024), and Auto-Guidance (Karras et al., 2024). For APG, we follow the recommendations from the paper and set $\gamma_{\text{APG}}=-0.5$, $\eta_{\text{APG}}=0.0$, $r_{\text{APG}}=5.0$. For CADS, we perform a grid search over $\tau_{1}^{\text{CADS}}\in[0.6,0.8],\tau_{2}^{\text{CADS}}\in[0.8,1.0],s^{\text{CADS}}\in[0.25,1.0],\psi^{\text{CADS}}=1.0$. Since SAG, PAG and SEG were developed specifically for the U-Net architecture, we adapt these method for diffusion transformers (Peebles and Xie, 2023) by applying the method in the attention layers of the middle blocks of the transformer; we refer to these methods with an asterisk superscript. Additional details on the choice of layers are provided in Section˜D.4. For AutoGuidance, we follow the recommendations from the paper and use an earlier checkpoint of the same model, at approximately $1/16$-th of the training, as the weaker model. To ensure a fair comparison, we select the best performing guidance strength for baselines as well as our method using the rank-scoring algorithm detailed in Section˜A.2. Note that this is different from reporting the optimal score achieved for each metric, which might not correspond to any particular run because of inherent trade-offs between the different facets of the generations.
@@ 5.2 Main experimental results
Figure 2: Pareto fronts on consistency-diversity-quality for text-to-image generation. Comparing ERG + APG (dots) with APG (crosses). We sweep over different guidance scales (each marked with a different color), and hyper-parameters $\alpha,\gamma,\tau$ for ERG. Dashed lines trace the Pareto fronts for each plot. We measure consistency with VQAScore, quality with density and diversity with coverage.
Text-to-image generation.  In Table˜1, we compare ERG with recent state-of-the-art guidance mechanisms on the text-to-image generation task. ERG demonstrates excellent performance, outperforming baselines such as CFG, SAG, PAG and SEG* in most metrics. Specifically, ERG considerably boosts image quality as reported for Density (e.g., +22 points when compared to CFG). Additionally, ERG achieves the highest consistency scores: +0.4+0.4 points in CLIPScore and +3.8+3.8 in VQAScore when compared to standard CFG. Moreover, ERG combined with APG achieves the overall best diversity, measured by FID and Coverage, and also improves Density over prior methods. These results suggest that our ERG approach is successfully able to boost the generation quality of the model in all three facets of generation (quality, diversity, and consistency). Moreover, the number of function evaluations (NFE) required for our approach is only two per inference step, which is comparable to other methods except SEG which requires three.
Following Astolfi et al. (2024), we plot Pareto fronts for quality (measured by Density), diversity (Coverage), and consistency (VQAScore) metrics in Figure˜2 when using ERG + APG. We find all the points belonging to the Pareto fronts correspond to ERG + APG, which improves in all three facets of the generations w.r.t. APG, and provides significant boosts for quality and consistency. This can also be observed in the qualitative comparison between APG and ERG + APG in Figure˜13 and Figure˜14 in the supplementary material. Similarly, in Figure˜1 the images sampled using ERG show better visual quality than those sampled with CFG.
@@
Figure 3: Unconditional generation results. Compared to not using guidance (top), our ERG generates more realistic and detailed images and more coherent structure (bottom). Images obtained from T2I model at $512$ with empty prompt as input. Samples in each column use the same seed.
Table 2: Unconditional generation. Comparing ERG to other approaches compatible with unconditional sampling.
FID $(\downarrow)$ Density $(\uparrow)$ Coverage $(\uparrow)$
No guidance $101.50$ $8.99$ $3.63$
SAG (Hong et al., 2023) $39.25$ $46.22$ $30.70$
PAG* (Ahn et al., 2024) $41.50$ $45.65$ $30.51$
SEG* (Hong, 2024) $37.75$ $55.56$ $34.79$
AutoGuidance (Karras et al., 2024) $39.50$ $48.26$ $34.71$
ERG (ours) $\bf 36.25$} $\bf 55.84$} $\bf 51.59$}
Table 3: Class-conditional generation. Comparison of ERG with other guidance methods for models trained for 256 and 512 resolution.
Res. FID $(\downarrow)$ Density $(\uparrow)$ Coverage $(\uparrow)$
CFG (Ho and Salimans, 2021) $256$ $\bf 3.67$} $127.03$ $85.81$
PAG* (Ahn et al., 2024) $5.31$ $111.94$ $81.04$
SAG (Hong et al., 2023) $3.78$ $131.89$ $86.35$
SEG* (Hong, 2024) $6.15$ $132.22$ $84.51$
ERG (ours) $\bf 3.67$} $\bf 141.96$} $\bf 86.72$}
CFG (Ho and Salimans, 2021) $512$ $5.65$ $146.97$ $\bf 86.70$}
PAG* (Ahn et al., 2024) $4.65$ $134.49$ $86.50$
SAG (Hong et al., 2023) $4.81$ $120.09$ $83.91$
SEG* (Hong, 2024) $6.59$ $160.11$ $81.85$
ERG (ours) $\bf 4.56$} $\bf 163.63$} $86.13$
Unconditional generation.  In the unconditional model sampling experiment, we compare I-ERG with sampling without guidance and using applicable methods: AutoGuidance, SEG*, PAG* and SAG*. We provide quantitative evaluation results in Table˜3, where we find that ERG outperforms all other methods, with significant boosts in FID, Density, and Coverage over other methods. The qualitative examples in Figure˜3 clearly exhibit artifacts in terms of structural coherence in all the objects present in the generated images when not using guidance, which disappear when using ERG.
Class-conditional generation.  For class-conditional generation, we observe similar trends as those seen for text-to-image and unconditional sampling in Table˜3. In particular, at 256 resolution, we find improvements in Density and Coverage. FID remains similar to CFG but is better compared to all other methods tested. At 512 resolution, ERG is best across all metrics, except for coverage where ERG is slightly behind CFG and SEG*.
Figure 4: Temperature rescaling in the conditioning for text-to-image generation (C-ERG). We vary the text encoder’s attention temperatures $\tau_{c}$. Each curve corresponds to a different guidance strength $w$. The left-most point on each curve represents the result for standard CFG.